Grade 11Physical Chemistry

If Z is a compressibility factr vander walls equation at liw pressure can be written as

keerthe sre
9 Years agoGrade 11

1 Answer

KUNAL PATEL
9 Years ago
VANDER WAAL EQUATION AFTER PRESSURE CORRECTION AND VOLUME EQUATION CAN BE GIVEN AS
(P+an2/V2)(V-nb)=nRT
At low pressures: Since V is large and therefore ‘b’ is negligible compared to V, the van der Waals equation for one mole of a gas is reduced to:
(P+a/V2)V=RT
PV+a/V=RT
PV=RT-a/V
DIVIDE WHOLE EQUATION BY RT.
PV/RT=1-a/VRT
AND AS WE KNOW THAT PV/RT=Z SO REPLACE PV/RT BY Z IN ABOVE EQUATION.
Z=1-a/VRT
AT VERY LOW PRESSURE:-Since V is very large and therefore both ‘b’ and ‘a/V2’ values become negligible. Hence the van der Waals equation for one mole of gas is reduced to: 
PV=RT
Therefore, at very low pressures all the gases obey the ideal gas equation.